A rectangular room has an area of 400 square foot. The length is 7 feet less than twice the width. How do you find the dimensions of the room?
Answer #1

400=w×(2w7). Length is 25 ft and width is 16 ft.

Explanation:

You can put an equation to solve the problem:

A=w×l

where l is 2w7

Now you can solve the question.

400=2w27w

w will be 16 feet.

I used trial and error method to find w. If you want, you can use quadratic formula to solve the question. If w is 16 ft,

400=(2×162)(7×16)

400=512112

It satisfies the equation.

Length is 25 ft and width is 16 ft.


answered by: sptutor